5. What is the 5th subnet of 172.16.0.0/21
Answer #1

Answer

Answer:
---------
172.16.32.0

Explanation:
-------------
172.16.0.0/21
Let's convert all octets to binary separately
Converting 172 to binary
Divide 172 successively by 2 until the quotient is 0
   > 172/2 = 86, remainder is 0
   > 86/2 = 43, remainder is 0
   > 43/2 = 21, remainder is 1
   > 21/2 = 10, remainder is 1
   > 10/2 = 5, remainder is 0
   > 5/2 = 2, remainder is 1
   > 2/2 = 1, remainder is 0
   > 1/2 = 0, remainder is 1
Read remainders from the bottom to top as 10101100
So, 172 of decimal is 10101100 in binary
172 in binary is 10101100

Converting 16 to binary
Divide 16 successively by 2 until the quotient is 0
   > 16/2 = 8, remainder is 0
   > 8/2 = 4, remainder is 0
   > 4/2 = 2, remainder is 0
   > 2/2 = 1, remainder is 0
   > 1/2 = 0, remainder is 1
Read remainders from the bottom to top as 10000
So, 16 of decimal is 10000 in binary
16 in binary is 00010000

Converting 0 to binary
Divide 0 successively by 2 until the quotient is 0
Read remainders from the bottom to top as
So, 0 of decimal is  in binary
0 in binary is 00000000

Converting 0 to binary
Divide 0 successively by 2 until the quotient is 0
Read remainders from the bottom to top as
So, 0 of decimal is  in binary
0 in binary is 00000000

=================================================================================
||    172.16.0.0 in binary notation is 10101100.00010000.00000000.00000000    ||
=================================================================================

10101100.00010000.00000000.00000000
/21 indicates that we are using next 5 bits for subnetting
first 5 subnets in the network are
10101100.00010000.00000000.00000000
10101100.00010000.00001000.00000000
10101100.00010000.00010000.00000000
10101100.00010000.00011000.00000000
10101100.00010000.00100000.00000000

5th subnet is 10101100.00010000.00100000.00000000
let's convert 3rd octet to decimal
=> 00100000
=> 0x2^7+0x2^6+1x2^5+0x2^4+0x2^3+0x2^2+0x2^1+0x2^0
=> 0x128+0x64+1x32+0x16+0x8+0x4+0x2+0x1
=> 0+0+32+0+0+0+0+0
=> 32
in decimal it is 172.16.32.0

5th subnet in the network is 172.16.32.0